The density function of coded measurements of pitch diameter of threads of a fitting is 𝑓(π‘₯) = { 2π‘₯ (π΅βˆ’1) (1 + π‘₯ 𝑩) 0 < π‘₯ < 1 0 π‘’π‘™π‘ π‘’π‘€β„Žπ‘’π‘Ÿπ‘’ Predict the probability of threads of a fitting is i. At least 0.5 ii. Anywhere 0.2 to 0.8


1
Expert's answer
2021-08-31T23:55:43-0400

Part i

P (threads of the fitting is at least 0.5)

"= P(X>0.5)\\
=\int_{0.5}^1f(x)dx=\int_{0.5}^1(\frac{2x^{B-1}}{1+x^B})dx"

But we can consider B as 6

"\int_{0.5}^1(\frac{2x^{6-1}}{1+x^6})dx\\
Let \space t= x^3 \space then \space dt = 3x^2dx \implies \frac{dt}{3}=x^2dx\\
\implies P(x>0.5) = \int_{0.5^3}^1 \frac{2t}{1+t^2}\frac{dt}{3}\\
= \frac{1}{3}\int_{0.5}^1 \frac{2t}{1+t^2}dt\\
= \frac{1}{3}[ \log (1+t^2)]_{0.5}^1\\
=\frac{1}{3} \log (\frac{16}{9})\\"


Part ii

"= P(0.2<X<0.8)\\
=\int_{0.2}^{0.8}f(x)dx=\int_{0.2}^{0.8}(\frac{2x^{B-1}}{1+x^B})dx"

But we can consider B as 6

"\int_{0.2}^{0.8}(\frac{2x^{6-1}}{1+x^6})dx\\
Let \space t= x^3 \space then \space dt = 3x^2dx \implies \frac{dt}{3}=x^2dx\\
\implies P(0.2<x<0.8) = \int_{0.2}^{0.8} \frac{2t}{1+t^2}\frac{dt}{3}\\
= \frac{1}{3}\int_{0.2}^{0.8} \frac{2t}{1+t^2}dt\\
= \frac{1}{3}[ \log (1+t^2)]_{0.2}^{0.8}\\
=\frac{1}{3} \log (\frac{3}{2})\\"